### v2

@@ 4.4 Optimistic Mirror Descent and Mirror Prox
$\mathbf{y}_{t}=\mathcal{P}_{\mathbf{x}_{t-1}}(\eta\nabla f(\mathbf{y}_{t-1})),$ (15)
$\mathbf{x}_{t}=\mathcal{P}_{\mathbf{x}_{t-1}}(\eta\nabla f(\mathbf{y}_{t})),$
@@
Theorem 10.
Under Assumptions 1, 2 and 3, if $0<\eta\leq 1/(3L)$, then the gradients along the trajectory generated by (15) satisfy $\max\{\left\|\nabla f(\mathbf{y}_{t})\right\|_{*},\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\}\leq G,\forall t\in\mathbb{N}$, and the convergence rate is given by
$f(\overline{\mathbf{y}}_{T})-f^{*}\leq\frac{D^{2}}{\eta T},$ (16)
where $\overline{\mathbf{y}}_{T}=\sum_{t=1}^{T}\mathbf{y}_{t}/T$.
@@
$\mathbf{y}_{t}=\mathcal{P}_{\mathbf{x}_{t-1}}(\eta\nabla f(\mathbf{x}_{t-1})),$ (17)
$\mathbf{x}_{t}=\mathcal{P}_{\mathbf{x}_{t-1}}(\eta\nabla f(\mathbf{y}_{t})).$
@@
The only difference between (15) and (17) is that the latter does not reuse gradient information, which results in $2$ gradient queries per iteration. For a comprehensive overview and comparison between these two algorithms, one may refer to Azizian et al. (2021); Cai et al. (2022b). The same $O(1/T)$ convergence rate is also obtained, as shown in the theorem below.
Theorem 11.
@@
Under Assumptions 1, 2 and 3, if $0<\eta\leq 1/(2L)$, then the gradients along the trajectory generated by (17) satisfy $\max\{\left\|\nabla f(\mathbf{y}_{t})\right\|_{*},\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\}\leq G,\forall t\in\mathbb{N}$, and the convergence rate is given by
$f(\overline{\mathbf{y}}_{T})-f^{*}\leq\frac{D^{2}}{\eta T},$ (18)
where $\overline{\mathbf{y}}_{T}=\sum_{t=1}^{T}\mathbf{y}_{t}/T$.
@@
The high-level idea behind Theorems 10 and 11 aligns with that in Section 4.3, i.e., bounding the dual norms of the gradients by analyzing suboptimality gaps. However, the implication from the last-iterate convergence in Theorems 8 and 9 to the bounded suboptimality gap does not apply for optimistic mirror descent and mirror prox, as elaborated below.
Hardness results on the last-iterate
@@
Different from the algorithms in Section 4.3 which enjoys a non-asymptotic last-iterate convergence under $L$-smoothness, the last-iterate of (15) as well as (17) had been only known to converge asymptotically, as shown by Popov (1980); Hsieh et al. (2019) for (15); Korpelevich (1976); Facchinei and Pang (2007) for (17). Even in the Euclidean setting, the finite-time last-iterate convergence is very challenging (Golowich et al., 2020; Wei et al., 2021; Gorbunov et al., 2022) and requires complicated analysis as well as intricate techniques (Cai et al., 2022a), e.g., computer-aided proofs based on sum-of-squares programming (Nesterov, 2000; Parrilo, 2003).
Circumvent the last-iterate analysis
@@
Recall the discussions in Section 4.2 that it suffices to ensure the boundness of the suboptimality gap, which can be implied by the last-iterate convergence. Since the latter one is a stronger statement, we can bypass the last-iterate convergence analysis, and control the suboptimality gap directly through careful analysis. Therefore, the cumbersome procedures in Cai et al. (2022a) can be avoided. Below, we briefly outline our approach.
Solution
@@
For the mirror prox method, we establish the descent property for both sequences (Lemma 25), i.e., $\max\left\{f(\mathbf{x}_{t}),f(\mathbf{y}_{t})\right\}\leq f(\mathbf{x}_{t-1})$, which guarantees the bounded suboptimality gap. For the optimistic mirror descent method, we track the stability term $\left\|\mathbf{y}_{t}-\mathbf{y}_{t-1}\right\|$ and show that it grows no faster than the geometric series (Lemma 23). The upper bound on stability terms suggests that the optimization trajectory is effectively stabilized, ultimately yielding $\max\left\{f(\mathbf{x}_{t}),f(\mathbf{y}_{t})\right\}\leq f(\mathbf{x}_{0})$. Consequently, we bound both the suboptimality gap and the gradient.
@@ 5 Stochastic Convex Optimization
In this section, we go one step further to study stochastic convex optimization (SCO) under $\ell*$-smoothness. Section 5.1 introduces the setup of SCO with a new noise assumption, while Section 5.2 gives the main result, providing a high-probability anytime convergence guarantee.
5.1 Setup
@@
where $\mathbf{g}_{t}$ is an estimate of the true gradient $\nabla f(\mathbf{x}_{t})$. Let $\bm{\epsilon}_{t}:=\mathbf{g}_{t}-\nabla f(\mathbf{x}_{t})$ denote the noise. We make the following assumption on $\bm{\epsilon}_{t}$, which is helpful for the high probability convergence.
Assumption 4
For all $t\in\mathbb{N}$, the stochastic gradients are unbiased:
$\mathbb{E}_{t-1}[\bm{\epsilon}_{t}]=0,$ (20)
@@
where the expectation $\mathbb{E}_{t-1}$ is conditioned on the past stochasticity $\left\{\bm{\epsilon}_{s}\right\}_{s=0}^{t-1}$. The noise satisfies the generalized bounded condition w.r.t. a non-decreasing polynomial function $\sigma:\mathbb{R}_{+}\mapsto\mathbb{R}_{++}$:

### v3

@@ 3.4 Optimistic Mirror Descent and Mirror Prox
$\mathbf{y}_{t}=\mathcal{P}_{\mathbf{x}_{t-1}}(\eta\nabla f(\mathbf{y}_{t-1})),\ \mathbf{x}_{t}=\mathcal{P}_{\mathbf{x}_{t-1}}(\eta\nabla f(\mathbf{y}_{t})),$ (10)
@@
Theorem 3.9.
Under Assumptions 2.5 and 3.3, if $0<\eta\leq 1/(3L)$, then the gradients along the trajectory generated by (10) satisfy $\max\{\left\|\nabla f(\mathbf{y}_{t})\right\|_{*},\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\}\leq G,\forall t\in\mathbb{N}$, and the convergence rate is given by
[…]
@@
$\mathbf{y}_{t}=\mathcal{P}_{\mathbf{x}_{t-1}}(\eta\nabla f(\mathbf{x}_{t-1})),\ \mathbf{x}_{t}=\mathcal{P}_{\mathbf{x}_{t-1}}(\eta\nabla f(\mathbf{y}_{t})).$ (12)
@@
The only difference between (10) and (12) is that the latter does not reuse gradient information, which results in $2$ gradient queries per iteration. For a comprehensive overview and comparison between these two algorithms, one may refer to Azizian et al. (2021); Cai et al. (2022a). The same $O(1/T)$ convergence rate is also obtained, as shown in the theorem below.
Theorem 3.10.
@@
Under Assumptions 2.5 and 3.3, if $0<\eta\leq 1/(2L)$, then the gradients along the trajectory generated by (12) satisfy $\max\{\left\|\nabla f(\mathbf{y}_{t})\right\|_{*},\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\}\leq G,\forall t\in\mathbb{N}$, and the convergence rate is given by
[…]
@@
The high-level idea behind Equations 11 and 13 aligns with that in Section 3.3, i.e., bounding the dual norms of the gradients by analyzing suboptimality gaps. However, the implication from the last-iterate convergence in Theorems 3.5 and 3.7 to the bounded suboptimality gap does not apply for optimistic mirror descent and mirror prox, as elaborated below.
@@
Hardness results on the last-iterate   Different from the algorithms in Section 3.3 which enjoys a non-asymptotic last-iterate convergence under $L$-smoothness, the last-iterate of (10) as well as (12) had been only known to converge asymptotically, as shown by Popov (1980); Hsieh et al. (2019) for (10); Korpelevich (1976); Facchinei and Pang (2007) for (12). Even in the Euclidean setting, the finite-time last-iterate convergence is very challenging (Golowich et al., 2020; Wei et al., 2021; Gorbunov et al., 2022) and requires complicated analysis as well as intricate techniques (Cai et al., 2022b), e.g., computer-aided proofs based on sum-of-squares programming (Nesterov, 2000; Parrilo, 2003).
@@
Circumvent the last-iterate analysis   Recall the discussions in Section 3.2 that it suffices to ensure the boundness of the suboptimality gap, which can be implied by the last-iterate convergence. Since the latter one is a stronger statement, we can bypass the last-iterate convergence analysis, and control the suboptimality gap directly through careful analysis. Therefore, the cumbersome procedures in Cai et al. (2022b) can be avoided. Below, we briefly outline our approach.
@@
Solution   For the mirror prox method, we establish the descent property for both sequences (Lemma H.2), i.e., $\max\left\{f(\mathbf{x}_{t}),f(\mathbf{y}_{t})\right\}\leq f(\mathbf{x}_{t-1})$, which guarantees the bounded suboptimality gap. For the optimistic mirror descent method, we track the stability term $\left\|\mathbf{y}_{t}-\mathbf{y}_{t-1}\right\|$ and show that it grows no faster than the geometric series (Lemma G.2). The upper bound on stability terms suggests that the optimization trajectory is effectively stabilized, ultimately yielding $\max\left\{f(\mathbf{x}_{t}),f(\mathbf{y}_{t})\right\}\leq f(\mathbf{x}_{0})$. Consequently, we bound both the suboptimality gap and the gradient.
@@ 4 Stochastic Convex Optimization
In this section, we extend from deterministic optimization to stochastic convex optimization.
4.1 Theoretical Results
@@
where $\mathbf{g}_{t}$ is an noisy estimate of the true gradient $\nabla f(\mathbf{x}_{t})$. Let $\boldsymbol{\epsilon}_{t}:=\mathbf{g}_{t}-\nabla f(\mathbf{x}_{t})$ denote the noise. We make the following sub-Gaussian assumption on $\boldsymbol{\epsilon}_{t}$, which is common for the high probability convergence (Juditsky et al., 2011; Vershynin, 2018; Zhang et al., 2023; Liu and Zhou, 2024).
Assumption 4.1.
@@
For all $t\in\mathbb{N}$, the stochastic gradients are unbiased: $\mathbb{E}_{t-1}[\boldsymbol{\epsilon}_{t}]=0$, where the expectation $\mathbb{E}_{t-1}$ is conditioned on the past stochasticity $\left\{\boldsymbol{\epsilon}_{s}\right\}_{s=0}^{t-1}$. The noise level further satisfies:
$\mathbb{E}_{t-1}\left[\exp\left(\lambda\left\|\boldsymbol{\epsilon}_{t}\right\|_{*}^{2}\right)\right]\leq\exp\left(\lambda\sigma^{2}\right),\forall\lambda\in[0,\sigma^{-2}].$
Now, we formally introduce the convergence of SMD, where the analysis is postponed to Appendix I.
Theorem 4.2.
Under Assumptions 2.5, 3.3 and 4.1, for any $0<\delta<1$, define
$G:=\sup\left\{\alpha\in\mathbb{R}_{+}|\alpha^{2}\leq 2\ell(2\alpha)\left(f(\mathbf{x}_{0})-f^{*}+\frac{64\sigma}{\sqrt{\delta}}\right)\right\},$
$\eta=\min\left\{\frac{64}{G},\sqrt{\frac{B(\mathbf{x}_{*},\mathbf{x}_{0})}{\sigma^{2}\log\left(\frac{1}{\delta}\right)\log(T)}}\right\},\quad L:=\ell(2G).$
Set $\eta_{t}=\min\left\{\frac{1}{2L},\frac{\eta}{\sqrt{T}}\right\}$, with probability at least $1-\delta$:
$f(\mathbf{x}_{T})-f^{*}$
$\leq$ $O\left(\frac{LB(\mathbf{x}_{*},\mathbf{x}_{0})}{T}+\sigma\sqrt{\frac{B(\mathbf{x}_{*},\mathbf{x}_{0})\log\left(\frac{1}{\delta}\right)\log(T)}{T}}\right).$
Moreover, $\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\leq G<+\infty$ holds for all $0\leq t\leq T-1$ with probability at least $1-\delta$.
The above theorem matches the state-of-the-art last-iterate convergence rate in Liu and Zhou (2024). Notably, our result is valid under $\ell*$-smooth function class, which is more general than the $(L,M)$-function class [^3] Liu and Zhou (2024, Assumption 3) state that $f$ is $(L,M)$-smooth such that $f(\mathbf{x})\leq f(\mathbf{y})+\langle\mathbf{g}_{\mathbf{y}},\mathbf{x}-\mathbf{y}\rangle+\frac{L}{2}\left\|\mathbf{x}-\mathbf{y}\right\|^{2}+M\left\|\mathbf{x}-\mathbf{y}\right\|,\forall\mathbf{x},\mathbf{y}\in\mathcal{X},\mathbf{g}_{\mathbf{y}}\in\partial f(\mathbf{y})$. in Liu and Zhou (2024). More importantly, our analysis is adaptive to noise, i.e., when $\sigma=0$ degenerates to deterministic optimization, the convergence in Theorem 4.2 encompasses that in Theorem 3.5 (deterministic mirror descent). Our theory answers affirmatively to the application of SMD to train real-world ML models, which is typically non-smooth and may satisfy our $\ell*$-smoothness condition (see our nanoGPT experiments in Section 2.3 as well as Appendix B).
Apart from the aforementioned convergence under sub-Gaussian noise, we also propose a new generalized bounded noise model (Assumption J.1) and establish a time-uniform convergence rate of $\widetilde{O}(1/\sqrt{t})$ (Theorem J.3). Due to space limitations, we defer this part to Appendix J.
4.2 Proof Sketch
Similar to the deterministic case, there are two major challenges underlying the convergence analysis for generalized smooth functions: (i) ensure in each step $t$, the iterates $\mathbf{x}_{t+1},\mathbf{x}_{t}$ are close enough to exploit the effective smooth property in Lemma 2.7; and (ii) ensure the local curvatures $\ell(2\left\|\nabla f(\mathbf{x}_{t})\right\|_{*})$ to be bounded uniformly by an absolute constant. For the second challenge, we can achieve this goal by bounding gradient norms $\left\{\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\right\}_{t=0}^{T-1}$, sharing a similar spirit as in Section 3. For the first challenge, the deterministic case begins with $\left\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\|\leq\eta\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}$. Since we have already controlled $\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}$ in the second challenge, the first challenge has been resolved simultaneously if the learning rate $\eta$ is set appropriately. However, for SCO, we write
$\left\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\|\leq\eta_{t+1}\left\|\mathbf{g}_{t}\right\|_{*}\leq\eta_{t+1}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}}+\eta_{t+1}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\left\|\boldsymbol{\epsilon}_{t}\right\|_{*}}.$
The gradient norm term $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}$} is effectively managed due to challenge (ii). For the noise term $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\left\|\boldsymbol{\epsilon}_{t}\right\|_{*}$}, we seek help from Assumption 4.1, which has good properties. We show that throughout $T$ iterations, the probability of “bad” noise estimates (with large noise norms $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\left\|\boldsymbol{\epsilon}_{t}\right\|_{*}$}) under Assumption 4.1 can be relatively low.
How do we manage $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}$} and $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\left\|\boldsymbol{\epsilon}_{t}\right\|_{*}$}?   We define the following events
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}A_{t}}:=\left\{f(\mathbf{x}_{t})-f^{*}\leq F\right\},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}B_{t}}:=\left\{\left\|\boldsymbol{\epsilon}_{t}\right\|_{*}\leq\frac{G}{2\eta L}\right\},$
and show that $\cup_{t=0}^{T-1}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}A_{t}}$ and $\cup_{t=0}^{T-1}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}B_{t}}$ happen simultaneously with high probability along the optimization trajectory. For $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}A_{t}$}, it is achieved by carefully analyzing the last-iterate behavior of (14): $f(\mathbf{x}_{t})-f(\mathbf{x})\leq b_{t}(B(\mathbf{x},\mathbf{x}_{0}))$, where $b_{t}$ denotes the bound depending on $B(\mathbf{x},\mathbf{x}_{0})$ for step $t$. We take the reference point $\mathbf{x}=\mathbf{x}_{0}$ to enable cancellation of terms, and obtain $f(\mathbf{x}_{t})-f(\mathbf{x}_{0})\leq b_{t}(0)$. By showing that $b_{t}$ is non-increasing in $t$, we are able to bound $f(\mathbf{x}_{t})-f^{*}\leq f(\mathbf{x}_{0})-f^{*}+\max_{t}b_{t}=O(1)$. For $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}B_{t}$}, we use Chebyshev’s inequality and take a union bound over the time horizon. The most subtle point underlying the analysis is that we have to perform a “chain-of-event” analysis, rather than taking union bounds brute-force. This is because when we analyze the failure probability of $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}A_{t}$}, previous events must happen, otherwise we are unable to exploit generalized smoothness. Such “chain-like” characteristics requires taking union bounds based on mathematical induction.
5 Non-convex Optimization
In this section, we extend $\ell*$-smoothness to non-convex composite optimization (Ghadimi et al., 2016; Lan, 2020), which can be formulated as
$F^{*}:=\min_{\mathbf{x}\in\mathcal{X}}\left\{F(\mathbf{x}):=f(\mathbf{x})+\phi(\mathbf{x})\right\}$ (15)
where $f\in\mathcal{F}_{\ell}(\left\|\cdot\right\|)$ and is possibly non-convex; $\phi:\mathcal{X}\mapsto\mathbb{R}_{+}$ is a simple convex regularizer, but possibly non-smooth (e.g., $\phi(\mathbf{x})\equiv 0$ or $\phi(\mathbf{x})=\left\|\mathbf{x}\right\|_{1}$). (15) can model a wide range of problems in ML since the objective $F$ is neither convex nor smooth. To optimize (15), the de facto method is composite mirror descent (CMD) (Duchi et al., 2010; Ghadimi et al., 2016; Lei and Tang, 2018; Lan, 2020; Wang et al., 2024b):
$\mathbf{x}_{t+1}=\operatorname*{argmin}_{\mathbf{x}\in\mathcal{X}}\left\{\langle\nabla f(\mathbf{x}_{t}),\mathbf{x}\rangle+\frac{B(\mathbf{x},\mathbf{x}_{t})}{\eta}+\phi(\mathbf{x})\right\},$
where we assume access to exact gradients. In the sequel, we study the convergence behavior of CMD under $\ell*$-smoothness and show that it still achieves the classic rate under our relaxed smoothness model.
To start, we define the following gradient mapping function (Nesterov and others, 2018) as the convergence criterion:
$\mathcal{G}_{t}:=\frac{\mathbf{x}_{t}-\mathbf{x}_{t+1}}{\eta},$ (16)
which is widely used in the context of mirror descent (Ghadimi et al., 2016; Lan, 2020; Huang et al., 2021). As pointed out by Lan (2020, Lemma 6.3), when the size of $\mathcal{G}_{t}$ vanishes, $\mathbf{x}_{t+1}$ approaches to a stationary point of (15). For more comprehensive illustrations, the readers may refer to Lan (2020), and references therein. Below, we formally present the theoretical guarantee, whose proof is located in Appendix K.
Theorem 5.1.
Under Assumptions 2.5 and 3.3, let
$G:=\sup\left\{\alpha\in\mathbb{R}_{+}|\alpha^{2}\leq 2\ell(2\alpha)\left(f(\mathbf{x}_{0})-f^{*}+\phi(\mathbf{x}_{0})\right)\right\},$
$L:=\ell(2G),\text{ and }0<\eta\leq\frac{1}{L}.$
Then, CMD enjoys the following convergence rate:
$\frac{1}{T}\sum_{t=0}^{t-1}\left\|\mathcal{G}_{t}\right\|^{2}\leq\frac{F(\mathbf{x}_{0})-F^{*}}{\eta T}.$ (17)
The convergence in (17) matches the one derived under the standard smoothness conditions (Ghadimi et al., 2016; Lan, 2020). This result provides strong theoretical support for the practical application of (composite) mirror descent, as modern neural networks are non-convex and non-smooth. Our result may also shed light on the application of CMD in LLM pretraining (Xie et al., 2023) and preference alignment (Munos et al., 2024; Zhang et al., 2025e, d; Wu et al., 2025), which has been empirically observed to exhibit generalized smooth properties (Crawshaw et al., 2022; Liu et al., 2025b; Riabinin et al., 2025).
6 Conclusion
In this paper, we propose a refined smoothness notion called $\ell*$-smoothness, which characterizes the norm of the Hessian by leveraging a general norm and its dual. This extension encompasses the current generalized smooth conditions while highlighting non-Euclidean problem geometries. We establish new convergence results of mirror descent, accelerated mirror descent, optimistic mirror descent, mirror prox, and stochastic mirror descent algorithms under this setting. Beyond convex optimization, we further delve into non-convex composite mirror descent and achieve tight convergence rates. Our theory provides insights into the practical usage of mirror descent, especially in pretraining and preference alignment of LLMs (Xie et al., 2023; Zhang et al., 2025e, d).
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Related Work
A.1 $(L_{0},L_{1})$-smoothness
$(L_{0},L_{1})$-smooth: non-convex case
The concept of $(L_{0},L_{1})$-smoothness is firstly proposed by Zhang et al. (2020b) based on empirical observations from LSTMs (Merity et al., 2018) and ResNets (He et al., 2016), which allows the function to have an affine-bounded Hessian norm. Under this new condition, they analyze gradient clipping (Mikolov and others, 2012; Pascanu et al., 2013) for deterministic non-convex optimization and derive a complexity of $O(\epsilon^{-2})$ [^4] In non-convex settings, the goal is to find $\epsilon$-stationary point $\mathbf{x}$ satisfying $\left\|\nabla f(\mathbf{x})\right\|_{2}\leq\epsilon$. In convex settings, we aim at $\epsilon$-suboptimal point $\mathbf{x}$ satisfying $f(\mathbf{x})-f^{*}\leq\epsilon$. The meaning of $\epsilon$ depends on the context., which matches the lower bound (Carmon et al., 2020) up to constant factors. They further extend to stochastic settings and provide an $O(\epsilon^{-4})$ complexity bound under the uniformly bounded noise assumption. Gaash et al. (2025) investigate the high-probability convergence of SGD with clipping, and obtain convergence rate that matches SGD up to polylogarithmic factors and additive terms. Yu et al. (2026) analyze both vector and matrix sign-based optimizers, namely SignSGD (Bernstein et al., 2018), Lion (Chen et al., 2023a), Muon (Jordan et al., 2024), and Muonlight (Liu et al., 2025a), under the generalized heavy-tailed noise conditions. They consider coordinate-wise $(L_{0},L_{1})$-smooth function class and obtain a complexity of $O(\epsilon^{-\frac{3p-2}{p-1}})$ for all methods, matching state-of-the-art.
$(L_{0},L_{1})$-smooth: convex case
Koloskova et al. (2023) show that gradient clipping has $O(\epsilon^{-1})$ complexity, matching the classic result in deterministic optimization. Takezawa et al. (2024) establish the same complexity bound for gradient descent with polyak stepsizes (Polyak, 1987). However, besides $(L_{0},L_{1})$-smoothness, these two works further impose an additional $L$-smooth assumption, where $L$ could be significantly larger than $L_{0}$ and $L_{1}$Gorbunov et al. (2025) address the limitation and recover their results without the extra $L$-smoothness assumption. Moreover, they study a variant of adaptive gradient descent (Malitsky and Mishchenko, 2020) and provide an $O(\epsilon^{-1})$ complexity result, albeit with worse constant terms compared to the original one. For acceleration schemes in convex optimization, they modify the method of similar triangles (MST) (Gasnikov and Nesterov, 2018) and prove an optimal complexity of $O(\epsilon^{-0.5})$Lobanov et al. (2024) provide a refined convergence analysis of gradient descent and achieve a linear speedup.
Other explorations
Ever since Zhang et al. (2020b) proposed $(L_{0},L_{1})$-smoothness condition, this generalized notion of smoothness has been flourishing in minimax optimization (Xian et al., 2024), bilevel optimization (Hao et al., 2024; Gong et al., 2024), multi-objective optimization (Zhang et al., 2025a), sign-based optimization (Crawshaw et al., 2022), zero-th order optimization (Lobanov and Gasnikov, 2025), distributionally robust optimization (Jin et al., 2021; Zhang et al., 2025b) and variational inequality (Vankov et al., 2024b, a). Numerous attempts have been made to refine existing algorithms under this weaker assumption, including variance reduction (Reisizadeh et al., 2025), clipping/normalized gradient (Zhang et al., 2020a; Qian et al., 2021; Zhao et al., 2021; Hübler et al., 2024; Yang et al., 2024), error feedback (Khirirat et al., 2024b) and trust region methods (Xie et al., 2024a). Notably, Faw et al. (2022, 2023); Wang et al. (2023b); Li et al. (2023b); Wang et al. (2024a); Zhang et al. (2025c) explore the convergence of AdaGrad (Duchi et al., 2011), RMSprop (Hinton et al., 2012), and Adam (Kingma and Ba, 2015) under generalized smoothness. Beyond the optimization community, it has also garnered significant attention in federated learning (Khirirat et al., 2024a; Demidovich et al., 2025), meta-learning (Chayti and Jaggi, 2024), and fine-tuning LLMs (Ma et al., 2025).
A.2 $\alpha$-symmetric Generalized Smoothness
An important generalization of $(L_{0},L_{1})$-smoothness is $\alpha$-symmetric $(L_{0},L_{1})$-smoothness proposed by Chen et al. (2023b), which introduces symmetry into the original formulation and also allows the dependency on the gradient norm to be polynomial with degree of $\alpha$. The formal definition is given as follows:
for some $L_{0},L_{1}\in\mathbb{R}_{+}$. For deterministic non-convex optimization, Chen et al. (2023b) establish the optimal complexity of $O(\epsilon^{-2})$ for a variant of normalized gradient descent (Nesterov, 1984; Cortés, 2006). They also show that the popular SPIDER algorithm (Fang et al., 2018) achieves the optimal $O(\epsilon^{-3})$ complexity in the stochastic setting. As pointed out by Chen et al. (2023b); Vankov et al. (2025), any twice-differentiable function satisfying (18) is also $(L_{0}^{\prime},L_{1}^{\prime})$-smooth for different constant factors. As an extension, Jiang et al. (2024) study sign-based finite-sum non-convex optimization and obtain an improved complexity over SignSVRG algorithm (Chzhen and Schechtman, 2023).
A.3 $\ell$-smoothness
A.4 Other Generalizations of Classic Smoothness
It is known that a function $f$ is $L$-smooth if LhfLh-f is convex (Nesterov and others, 2018) for $h(\cdot)=\left\|\cdot\right\|^{2}/2$. Advancements have been made to relax $h$ into an arbitrary Legendre function, leading to the development of a notion called relative smoothness (Birnbaum et al., 2011; Bauschke et al., 2017; Lu et al., 2018; Hanzely et al., 2021; Hanzely and Richtárik, 2021), which substitutes the canonical quadratic upper bound with $f(\mathbf{x})\leq f(\mathbf{y})+\langle\nabla f(\mathbf{y}),\mathbf{x}-\mathbf{y}\rangle+LB_{h}(\mathbf{x},\mathbf{y})$, where $B_{h}$ is the Bregman divergence associated with $h$. Since relative smoothness is now standard in the analysis of mirror descent, we provide a detailed comparison with our $\ell*$-smoothness in Section A.5.
More recently, Mishkin et al. (2024) propose directional smoothness, a measure of local gradient variation that preserves the global smoothness along specific directions. To address the imbalanced Hessian spectrum distribution in practice (Sagun et al., 2016; Pan et al., 2022)Liu et al. (2025b); Jiang et al. (2025) propose anisotropic smoothness and obtain improved bounds for adaptive gradient methods.
A.5 Comparison with Relative Smoothness
According to Lu et al. (2018), a function $f$ is $L$-smooth relative to $h$ iff LhfLh-f is convex, which is equivalent to $\nabla^{2}f(x)\preceq L\nabla^{2}h(x)$ when both functions are twice-differentiable on the interior of the domain. Notably, $h$ is not presumed to possess any special properties like strict or strong convexity. In the sequel, we compare this notion with $\ell*$-smoothness and highlight why ours is more practical.
Practical infeasibility
Relative smoothness constrains the Hessian spectrum via a reference function $h$. However, for any $f$, there exists infinitely many valid $(L,h)$ pairs satisfying the definition. Therefore, finding the most appropriate pair is essential under this framework. Unfortunately, in most cases, $(L,h)$ can only be determined through delicate design (requiring prior knowledge of $f$). In contrast, $\ell*$-smoothness bounds the matrix primal-dual norm of the Hessian via a link function of the gradient dual norm, i.e., $\left\|\nabla^{2}f(\mathbf{x})\right\|_{*,\cdot}:=\sup_{\mathbf{h}}\left\{\left\|\nabla^{2}f(\mathbf{x})\mathbf{h}\right\|_{*}/\left\|\mathbf{h}\right\|\right\}\leq\ell(\left\|\nabla f(\mathbf{x})\right\|_{*})$. This formulation does not involve Löwner partial order $\preceq$ and is generally much easier to verify, as it compresses the information of eigenvalues into a specific norm and connects it with gradients. Moreover, recent advances in ML suggest that $\ell*$-smoothness is satisfied by many practical objectives both theoretically and empirically (see references discussed earlier), whereas relative smoothness remains less prevalent in the ML community. To summarize, both notions extend the traditional smoothness from different perspectives: relative smoothness is more direct and strict, while generalized smoothness is more flexible and practical. Below, we provide concrete examples and further reasoning to support our claims.
Examples
Key limitations
  1. 1.
  2. 2.
Appendix B Experimental Details for LLMs
(a) Validation of layer-wise $(L^{0},L^{1})$-smoothness for the group of parameters from the transformer blocks of NanoGPT-124M along unScion training trajectories. The group norms are non-euclidean norms $\|\cdot\|_{(i)}=\sqrt{n_{i}/m_{i}}\|\cdot\|_{2\to 2}$.
All experiments for the NanoGPT model (Karpathy, 2022) are conducted using PyTorch [^5] PyTorch Documentation. Available at: https://pytorch.org/docs/stable/index.html with Distributed Data Parallel (DDP) [^6] Distributed Data Parallel (DDP) in PyTorch. Available at: https://pytorch.org/docs/stable/notes/ddp.html across 8 NVIDIA 4090 GPUs (24GB each). The experiments are based on an open-source codebase which can be found at https://github.com/artem-riabinin/Experiments-estimating-smoothness-for-NanoGPT-and-CNN/. Theoretical formulations and details for the codebase can be found at Riabinin et al. (2025, Appendix E).
We mainly aimed at empirically validating our generalized smoothness model under non-Euclidean norms. In the procedure of training nanoGPT on the Fineweb dataset (Penedo et al., 2024) with unScion (Riabinin et al., 2025), we plot the estimated trajectory smoothness
$\hat{L}_{i}[k]\coloneqq\frac{\|\nabla_{i}f_{\xi^{k+1}}(X^{k+1})-\nabla_{i}f_{\xi^{k}}(X^{k})\|_{(i)\star}}{\|X_{i}^{k+1}-X_{i}^{k}\|_{(i)}},$
and its approximation
$\hat{L}_{i}^{\text{approx}}[k]\coloneqq L_{i}^{0}+L_{i}^{1}\|\nabla_{i}f_{\xi^{k+1}}(X^{k+1})\|_{(i)\star}$
as functions of the iteration index $k$. We observed similar findings to Riabinin et al. (2025, Appendix E.3.1), where a there exihibits clear alignment between $\hat{L}_{i}[k]$ and $\hat{L}_{i}^{\text{approx}}[k]$, suggesting that our smoothness model holds approximately under non-Euclidean norms. Figure 5(a) shows more results for parameter groups at different transformer blocks of NanoGPT-124M.
Appendix C The Gain of Algorithmic Adaptivity
In this section, we present the omitted details for Section 2.2. The proof of Proposition 2.8 is essentially the discussions underlying Section C.1, which are inspired by Gorbunov et al. (2025). First, we present the omitted notations in the main text.
Omitted notations